# BasicAVSR: Arbitrary-Scale Video Super-Resolution via Image Priors and Enhanced Motion Compensation

### v1

@@ 1 Introduction
The evolutionary and developmental processes of our visual systems have presumably been shaped by continuous visual data [56]. Yet, how to acquire and represent a natural scene as a continuous signal remains wide open. This difficulty stems from two main factors. The first is the physical limitations of digital imaging devices, including sensor size and density, optical diffraction, lens quality, electrical noise, and processing power. The second is the inherent complexities of natural scenes, characterized by their wide and deep frequencies, which pose significant challenges for applying the Nyquist–Shannon sampling [44] and compressed sensing [14] theories to accurately reconstruct continuous signals from discrete samples. Consequently, natural scenes are predominantly represented as discrete pixel arrays, often with limited resolution.
@@
Super-resolution (SR) provides an effective means of enhancing the resolution of low-resolution (LR) images and videos [23, 50]. Early deep learning-based SR methods [13, 51, 34, 61] focus on fixed integer scaling factors (e.g., $\times 4$ and $\times 8$), each corresponding to an independent convolutional neural network (CNN). This limits their applicability in real-world scenarios, where varying scaling requirements are common. From the human vision perspective, users may want to continuously zoom in on images and videos to arbitrary scales using the two-finger pinch-zoom feature on mobile devices as a natural form of human-computer interaction. From the machine vision perspective, different applications (such as computer-aided diagnosis, remote sensing, and video surveillance) may require different scaling factors to zoom in on different levels of detail for optimal analysis and decision-making.
@@
Recently, arbitrary-scale image SR (AISR) has gained significant attention due to its capability of upsampling LR images to arbitrary high-resolution (HR) using a single model. Contemporary AISR methods can be categorized into four classes based on how arbitrary-scale upsampling is performed: interpolation-based methods [26, 1], learnable adaptive filter-based methods [22, 57, 59], implicit neural representation-based methods [10, 30, 9], and Gaussian splatting-based methods [21, 45]. These algorithms face several limitations, including quality degradation at high (and possibly integer) scales [22, 10, 57], high computational complexity [30, 9], and difficulty in generalizing across unseen scales and degradation models [22, 10, 30, 9], as well as temporal inconsistency in video SR.
@@
Compared to AISR, arbitrary-scale video SR (AVSR) is significantly more challenging due to the added time dimension. Existing AVSR methods [12, 11, 25] rely primarily on conditional neural radiance fields [40] as continuous signal representations. Due to the high computational demands during training and inference, only two adjacent frames are used for spatiotemporal modeling, which is bound to be suboptimal. Another method [31] adopts an upsampling kernel, thereby avoiding the drawbacks of conditional neural radiance fields. However, treating scale factors as priors can not provide guidance related to image content. The sliding-window-based bidirectional RNN utilized in this method has limitations in modeling long-sequence motions and is computationally inefficient. Moreover, current AVSR approaches rely on optical flows for inter-frame alignment; yet, imprecise optical flow estimation degrades model performance.
@@
In this work, we aim for AVSR with the goal of reproducing faithful spatial detail and maintaining coherent temporal consistency at low computational complexity. We describe a strong baseline, which we name BasicAVSR, by identifying and combining four variants of elementary building blocks [7, 2]: 1) multi-scale frequency priors, 2) a flow-guided propagation unit, 3) a second-order motion compensation unit, and 4) a hyper-upsampling unit. BasicAVSR is grounded in scale-space theory [35] in computer vision and image processing, which suggests that human perception and interpretation of real-world structures and textures are scale-dependent. As shown in Fig. 1, multi-scale frequency priors offer rich frequency-domain information, which is beneficial for restoring detailed textures and structures in VSR. The same image reveals distinct frequency-band spatial differences across different input resolutions. Accurately characterizing these multi-scale image frequency bands is highly valuable for AVSR. We extract frequency-domain priors adaptively from the Laplacian pyramid decomposition of each frame. These priors effectively distinguish structures and textures across scales and capture mid-level visual concepts tied to image layout [15, 62]. The flow-guided propagation unit captures long-term spatiotemporal dependencies from adjacent frames, thereby improving the temporal consistency of the video sequence. The second-order motion compensation unit refines the spatial alignment of adjacent frames and searches for regions with similar image content near the initially estimated motion offset, thereby achieving more accurate motion estimation. The hyper-upsampling unit trains a hyper-network [18] that takes scale-relevant parameters as input to generate content-independent upsampling kernels, enabling pre-computation to accelerate inference speed.
@@
This work is primarily presented as a conference paper [49], upon which this manuscript has made two major improvements, i.e., a more direct way of generating multi-scale priors is introduced to eliminate the need for extra pre-trained networks and a motion compensation mechanism is introduced to enhance the accuracy of video frames alignment. In addition, the method is extended to multiple application scenarios, including online, offline, and quasi-online settings. To sum up, the main contributions of this work include:
  • A strong baseline, BasicAVSR, that is a nontrivial combination of four variants of elementary building blocks in literature [7, 2],
@@
  • A comprehensive experimental demonstration, that BasicAVSR significantly surpasses competing methods in terms of SR quality on different test sets, generalization ability to unseen scales and degradation models, as well as inference speed.
@@ 2.1 Key Components of VSR
Kappeler et al[24] pioneered CNN-based approaches for VSR, emphasizing two key components: feature alignment and aggregation. Subsequent studies have focused on enhancing these components. EDVR [58] introduced pyramid deformable alignment and spatiotemporal attention for feature alignment and aggregation. BasicVSR [6] and BasicVSR++++ [7] employ an optical flow-based module to estimate motion correspondence between neighboring frames for feature alignment and a bidirectional propagation module to aggregate spatiotemporal information from previous and future frames, which set the VSR performance record at that time. RVRT [33] enhanced VSR performance by utilizing a recurrent video restoration Transformer with guided deformable attention albeit at the expense of substantially increased computational complexity. Additionally, VideoINR [12], MoTIF [11], and BF-STVSR [25] integrated VSR with video frame interpolation, which achieved limited success due to the ill-posedness of the task. In our work, we combine a flow-guided propagation unit and a second-oreder motion compensation unit to extract, align, and aggregate spatiotemporal features from adjacent frames, while keeping computational complexity manageable.
@@ 2.2 Upsampling Modules for AISR and AVSR
Compared to fixed-scale SR methods  [13, 51, 34, 61, 32], upsampling plays a more crucial role in AISR and AVSR. Besides direct interpolation-based upsampling [26, 1], learnable adaptive filter-based upsampling, implicit neural representation-based upsampling, and Gaussian splatting-based upsampling are commonly used. Meta-SR [22] was the pioneer in AISR, dynamically predicting the upsampling kernels using a single model. ArbSR [57] introduced a scale-aware upsampling layer compatible with fixed-scale SR methods. EQSR [59] proposed a bilateral encoding of both scale-aware and content-dependent features during upsampling. Inspired by the success of implicit neural representations in computer graphics [39, 46], this approach has also been applied to AISR and AVSR. For instance, LIIF [10] predicts the RGB values of HR pixels using the coordinates of LR pixels along with their neighboring features as inputs. LTE [30] captures more fine detail with a local texture estimator, and CLIT [9] enhances representation expressiveness with cross-scale attention and multi-scale reconstruction. OPE [53] introduced orthogonal position encoding for efficient upsampling. CiaoSR [4] proposed an attention-based weight ensemble algorithm for feature aggregation in a large receptive field. Recently, 2D Gaussian Splatting has shown great potential in image processing [21, 45]. Unlike traditional methods that treat pixels as discrete points, Gaussian splatting-based upsampling represents each pixel as a continuous Gaussian field. By rendering mutually stacked Gaussian fields, the encoded features are simultaneously refined and upsampled, which establishes long-range dependencies and enhances representation ability.
Existing AVSR methods [12, 11, 25] also use implicit neural representations but are constrained to modeling spatiotemporal relationships between only two adjacent frames due to the high computational costs involved. In contrast to the these approaches, SAVSR [31] introduces a dual-branch upsampling architecture. It adaptively adjusts network weights based on spatiotemporal features and upsampling scales. However, the computational process of both branches relies on input features, which means pre-calculation is not feasible in this architecture. The proposed BasicAVSR addresses this limitation by employing a lightweight hyper-upsampling unit to predict scale-aware and content-independent upsampling kernels, allowing for pre-computation to speed up inference.
@@ 2.3 Natural Scene Priors for SR
The history of SR, or more generally low-level vision, is closely tied to the development of natural scene priors. Commonly used priors in SR include the smoothness prior [5], sparsity prior [38], self-similarity prior [17], edge/gradient prior [19], deep architectural prior [55], temporal consistency prior [3], motion prior [54, 48], and perceptual prior [60]. In the subfield of AISR and AVSR, the scaling factor-based priors have exclusively been leveraged as adaptive convolution conditions [57, 16, 59, 31]. However, single scaling factors are limited in providing rich texture priors that vary with the video content. In this paper, we introduce multi-scale frequency priors. Using Laplacian pyramid decomposition, we effectively distinguish between high- and low-frequency information of images at varying locations and scales. We demonstrate its effectiveness in enhancing AVSR.
@@ 3.1 Multi-Scale Frequency Priors for AVSR
Accurately characterizing image structure and texture at multiple scales is crucial for the task of AVSR. Fortunately, the scale-space theory in computer vision and image processing [28, 35] provides an elegant theoretical framework for this purpose. Since different frequency bands of an image can present the structure and texture of the image, in this work, we adopt multi-scale frequency priors derived from Laplacian pyramid decomposition as an alternative to the VGG-based features used in our previous work [49]. This approach breaks down an image into multiple layers of different frequency bands. The resulting pyramid consists of several levels, each representing the image at a different scale and capturing specific frequency information. As illustrated in Fig. 1, different frequency bands provide explicit information about the spatial distribution of details and textures at various scales. By replacing the deep-learning-based features with frequency-domain information obtained through Laplacian pyramid decomposition, we achieve comparable performance while eliminating the need for pre-trained networks (please refer to the analysis in section 4.4 for details). This reduces both the parameters and inference time. Specifically, we upsample different frequency bands to match the input resolution and apply learnable weights for band-by-band weighting. This allows the network to adaptively adjust the importance of different frequency bands for diverse videos. Next, we concatenate the fused frequency maps with the current frame $\bm{x}_{i}$, which serves as the multi-scale frequency prior, denoted by $\bm{p}_{i}$. Inserting these frequency priors into the current model is straightforward: we replace all instances of $\bm{x}$ with $\bm{p}$ (except for the last residual connection which produces the HR video $\hat{\bm{y}}$).
@@ 3.2 Bidirectional Flow-Guided Propagation Unit
Given $\bm{x}=\{\bm{x}_{i}\}_{i=0}^{T}$, the bidirectional flow-guided propagation unit computes two sequence of hidden states $\{\bm{h}^{\rightarrow}_{i}\}_{i=0}^{T}$ and $\{\bm{h}^{\leftarrow}_{i}\}_{i=0}^{T}$ to capture long-term spatiotemporal dependencies of previous and future frames. It is worth emphasizing that we first fed the features with priors $\bm{p}$ into a ResNet with $N$ residual blocks for feature extraction, which serve as the original feature list $\bm{h}$. We take estimating the forward hidden states $\{\bm{h}^{\rightarrow}_{i}\}_{i=0}^{T}$ as an example. Initially, we estimate the optical flow between the current and previous frames [6]:
@@
where $\mathtt{flow}(\cdot)$ denotes a state-of-the-art optical flow estimator [47]. $\bm{f}_{i\rightarrow{i-1}}$ is then used to align the hidden state $\bm{h}_{i-1}$ backward:
@@
where $\mathtt{calign}(\cdot)$ denotes alignment based on motion compensation (discussed in the next subsection). Subsequently, the aligned previous hidden state $\bm{h}_{i-1\rightarrow i}$ and the current frame $\bm{x}_{i}$ are concatenated along the channel dimension and processed through a ResNet with $N$ residual blocks to compute $\bm{h}^{\rightarrow}_{i}$. The backward hidden state $\bm{h}^{\leftarrow}_{i}$ can be computed by reversing the input sequence and applying the aforementioned formulas. It also needs to be concatenated with the previously computed hidden state and fed into a ResNet with $N$ residual blocks. As prior studies [7] have demonstrated, multiple bidirectional information interaction refines features by propagating intermediate features forward and backward in an alternating manner over time. This allows information from different frames to be revisited for feature refinement. Hence, we employ a two-iteration bidirectional RNN propagation to compute bidirectional hidden states. Specifically, after obtaining forward and backward hidden states via the aforementioned operations, we repeat the process and substitute the original hidden states with the refined ones. To further enhance propagation robustness, we adopt second-order connections, which aggregate information from more spatiotemporal locations, improving the robustness and effectiveness of occluded and fine-detailed regions. More details are provided in the next section. The bidirectional flow-guided propagation unit allows the proposed BasicAVSR to incorporate long-term spatialtemporal context while being flow-aware.

### v2

@@ 1 Introduction
The evolutionary and developmental processes of our visual systems have presumably been shaped by continuous visual data [wandell1995foundations]. Yet, how to acquire and represent a natural scene as a continuous signal remains wide open. This difficulty stems from two main factors. The first is the physical limitations of digital imaging devices, including sensor size and density, optical diffraction, lens quality, electrical noise, and processing power. The second is the inherent complexities of natural scenes, characterized by their wide and deep frequencies, which pose significant challenges for applying the Nyquist–Shannon sampling [oppenheim1997signals] and compressed sensing [donoho2006compressed] theories to accurately reconstruct continuous signals from discrete samples. Consequently, natural scenes are predominantly represented as discrete pixel arrays, often with limited resolution.
@@
Super-resolution (SR) provides an effective means of enhancing the resolution of low-resolution (LR) images and videos [irani1991improving, shechtman2005space]. Early deep learning-based SR methods [dong2014learning, shi2016real, lim2017enhanced, zhang2018residual] focus on fixed integer scaling factors (e.g., $\times 4$ and $\times 8$), each corresponding to an independent convolutional neural network (CNN). This limits their applicability in real-world scenarios, where varying scaling requirements are common. From the human vision perspective, users may want to continuously zoom in on images and videos to arbitrary scales using the two-finger pinch-zoom feature on mobile devices as a natural form of human-computer interaction. From the machine vision perspective, different applications (such as computer-aided diagnosis, remote sensing, and video surveillance) may require different scaling factors to zoom in on different levels of detail for optimal analysis and decision-making.
@@
Recently, arbitrary-scale image SR (AISR) has gained significant attention due to its capability of upsampling LR images to arbitrary high-resolution (HR) using a single model. Contemporary AISR methods can be categorized into four classes based on how arbitrary-scale upsampling is performed: interpolation-based methods [kim2016accurate, behjati2021overnet], learnable adaptive filter-based methods [hu2019meta, wang2021learning, wang2023deep], implicit neural representation-based methods [chen2021learning, lee2022local, chen2023cascaded], and Gaussian splatting-based methods [hu2025gaussiansr, peng2025pixel]. These algorithms face several limitations, including quality degradation at high (and possibly integer) scales [hu2019meta, chen2021learning, wang2021learning], high computational complexity [lee2022local, chen2023cascaded], and difficulty in generalizing across unseen scales and degradation models [hu2019meta, chen2021learning, lee2022local, chen2023cascaded], as well as temporal inconsistency in video SR.
@@
Compared to AISR, arbitrary-scale video SR (AVSR) is significantly more challenging due to the added time dimension. Existing AVSR methods [chen2022videoinr, chen2023motif, kim2025bf] rely primarily on conditional neural radiance fields [mildenhall2020nerf] as continuous signal representations. Due to the high computational demands during training and inference, only two adjacent frames are used for spatiotemporal modeling, which is bound to be suboptimal. Another method [li2024savsr] adopts an upsampling kernel, thereby avoiding the drawbacks of conditional neural radiance fields. However, treating scale factors as priors can not provide guidance related to image content. The sliding-window-based bidirectional RNN utilized in this method has limitations in modeling long-sequence motions and is computationally inefficient. Moreover, current AVSR approaches rely on optical flows for inter-frame alignment; yet, imprecise optical flow estimation degrades model performance.
@@
In this work, we aim for AVSR with the goal of reproducing faithful spatial detail and maintaining coherent temporal consistency at low computational complexity. We describe a strong baseline, which we name BasicAVSR, by identifying and combining four variants of elementary building blocks [chan2022basicvsr++, burt1983laplacian]: 1) multi-scale frequency priors, 2) a flow-guided propagation unit, 3) a second-order motion compensation unit, and 4) a hyper-upsampling unit. BasicAVSR is grounded in scale-space theory [lindeberg2013scale] in computer vision and image processing, which suggests that human perception and interpretation of real-world structures and textures are scale-dependent. As shown in Fig. 1, multi-scale frequency priors offer rich frequency-domain information, which is beneficial for restoring detailed textures and structures in VSR. The same image reveals distinct frequency-band spatial differences across different input resolutions. Accurately characterizing these multi-scale image frequency bands is highly valuable for AVSR. We extract frequency-domain priors adaptively from the Laplacian pyramid decomposition of each frame. These priors effectively distinguish structures and textures across scales and capture mid-level visual concepts tied to image layout [fu2023dreamsim, zhou2023dichotomous]. The flow-guided propagation unit captures long-term spatiotemporal dependencies from adjacent frames, thereby improving the temporal consistency of the video sequence. The second-order motion compensation unit refines the spatial alignment of adjacent frames and searches for regions with similar image content near the initially estimated motion offset, thereby achieving more accurate motion estimation. The hyper-upsampling unit trains a hyper-network [ha2017hypernetworks] that takes scale-relevant parameters as input to generate content-independent upsampling kernels, enabling pre-computation to accelerate inference speed.
@@
This work is primarily presented as a conference paper [shang2024arbitrary], upon which this manuscript has made two major improvements, i.e., a more direct way of generating multi-scale priors is introduced to eliminate the need for extra pre-trained networks, and a motion compensation mechanism is introduced to enhance the accuracy of video frames alignment. In addition, the method is extended to multiple application scenarios, including online, offline, and quasi-online settings. To sum up, the main contributions of this work include:
  • A strong baseline, BasicAVSR, that is a nontrivial combination of four variants of elementary building blocks in literature [chan2022basicvsr++, burt1983laplacian],
@@
  • A comprehensive experiment that shows BasicAVSR significantly surpasses competing methods in terms of SR quality on different test sets, generalization ability to unseen scales and degradation models, as well as inference speed.
@@ 2.1 Key Components of VSR
Kappeler et al[kappeler2016video] pioneered CNN-based approaches for VSR, emphasizing two key components: feature alignment and aggregation. Subsequent studies have focused on enhancing these components. EDVR [wang2019edvr] introduced pyramid deformable alignment and spatiotemporal attention for feature alignment and aggregation. BasicVSR [chan2021basicvsr] and BasicVSR++++ [chan2022basicvsr++] employ an optical flow-based module to estimate motion correspondence between neighboring frames for feature alignment and a bidirectional propagation module to aggregate spatiotemporal information from previous and future frames, which set the VSR performance record at that time. RVRT [liang2022recurrent] enhanced VSR performance by utilizing a recurrent video restoration Transformer with guided deformable attention albeit at the expense of substantially increased computational complexity. Additionally, VideoINR [chen2022videoinr], MoTIF [chen2023motif], and BF-STVSR [kim2025bf] integrated VSR with video frame interpolation, which achieved limited success due to the ill-posedness of the task. In our work, we combine a flow-guided propagation unit and a second-order motion compensation unit to extract, align, and aggregate spatiotemporal features from adjacent frames, while keeping computational complexity manageable.
@@ 2.2 Upsampling Modules for AISR and AVSR
Compared to fixed-scale SR methods  [dong2014learning, shi2016real, lim2017enhanced, zhang2018residual, liang2021swinir], upsampling plays a more crucial role in AISR and AVSR. Besides direct interpolation-based upsampling [kim2016accurate, behjati2021overnet], learnable adaptive filter-based upsampling, implicit neural representation-based upsampling, and Gaussian splatting-based upsampling are commonly used. Meta-SR [hu2019meta] was the pioneer in AISR, dynamically predicting the upsampling kernels using a single model. ArbSR [wang2021learning] introduced a scale-aware upsampling layer compatible with fixed-scale SR methods. EQSR [wang2023deep] proposed a bilateral encoding of both scale-aware and content-dependent features during upsampling. Inspired by the success of implicit neural representations in computer graphics [michalkiewicz2019implicit, peng2020convolutional], this approach has also been applied to AISR and AVSR. For instance, LIIF [chen2021learning] predicts the RGB values of HR pixels using the coordinates of LR pixels along with their neighboring features as inputs. LTE [lee2022local] captures more fine detail with a local texture estimator, and CLIT [chen2023cascaded] enhances representation expressiveness with cross-scale attention and multi-scale reconstruction. OPE [song2023ope] introduced orthogonal position encoding for efficient upsampling. CiaoSR [cao2023ciaosr] proposed an attention-based weight ensemble algorithm for feature aggregation in a large receptive field. Recently, 2D Gaussian Splatting has shown great potential in image processing [hu2025gaussiansr, peng2025pixel]. Unlike traditional methods that treat pixels as discrete points, Gaussian splatting-based upsampling represents each pixel as a continuous Gaussian field. By rendering mutually stacked Gaussian fields, the encoded features are simultaneously refined and upsampled, which establishes long-range dependencies and enhances representation ability.
Existing AVSR methods [chen2022videoinr, chen2023motif, kim2025bf] also use implicit neural representations but are constrained to modeling spatiotemporal relationships between only two adjacent frames due to the high computational costs involved. In contrast to these approaches, SAVSR [li2024savsr] introduces a dual-branch upsampling architecture. It adaptively adjusts network weights based on spatiotemporal features and upsampling scales. However, the computational process of both branches relies on input features, which means pre-calculation is not feasible in this architecture. The proposed BasicAVSR addresses this limitation by employing a lightweight hyper-upsampling unit to predict scale-aware and content-independent upsampling kernels, allowing for pre-computation to speed up inference.
@@ 2.3 Natural Scene Priors for SR
The history of SR, or more general low-level vision, is closely tied to the development of natural scene priors. Commonly used priors in SR include the smoothness prior [chambolle2004algorithm], sparsity prior [mairal2014sparse], self-similarity prior [glasner2009super], edge/gradient prior [he2011single], deep architectural prior [ulyanov2018deep], temporal consistency prior [caballero2017real], motion prior [tao2017detail, shang2023joint], and perceptual prior [wang2018recovering]. In the subfield of AISR and AVSR, the scaling factor-based priors have exclusively been leveraged as adaptive convolution conditions [wang2021learning, fu2021residual, wang2023deep, li2024savsr]. However, single scaling factors are limited in providing rich texture priors that vary with the video content. In this paper, we introduce multi-scale frequency priors. Using Laplacian pyramid decomposition, we effectively distinguish between high- and low-frequency information of images at varying locations and scales. We demonstrate its effectiveness in enhancing AVSR.
@@ 3.1 Multi-Scale Frequency Priors for AVSR
Accurately characterizing image structure and texture at multiple scales is crucial for the task of AVSR. Fortunately, the scale-space theory in computer vision and image processing [koenderink1984structure, lindeberg2013scale] provides an elegant theoretical framework for this purpose. Since different frequency bands of an image can present the structure and texture of the image, in this work, we adopt multi-scale frequency priors derived from Laplacian pyramid decomposition as an alternative to the VGG-based features used in our previous work [shang2024arbitrary]. This approach breaks down an image into multiple layers of different frequency bands. The resulting pyramid consists of several levels, each representing the image at a different scale and capturing specific frequency information. As illustrated in Fig. 1, different frequency bands provide explicit information about the spatial distribution of details and textures at various scales. By replacing the deep-learning-based features with frequency-domain information obtained through Laplacian pyramid decomposition, we achieve comparable performance while eliminating the need for pre-trained networks (please refer to the analysis in section 4.4 for details). This reduces both the parameters and inference time. Specifically, we upsample different frequency bands to match the input resolution and apply learnable weights for band-by-band weighting. This allows the network to adaptively adjust the importance of different frequency bands for diverse videos. Next, we concatenate the fused frequency maps with the current frame $\bm{x}_{i}$, which serves as the multi-scale frequency prior, denoted by $\bm{p}_{i}$. Inserting these frequency priors into the current model is straightforward: we replace all instances of $\bm{x}$ with $\bm{p}$ (except for the last residual connection which produces the HR video $\hat{\bm{y}}$).
@@ 3.2 Bidirectional Flow-Guided Propagation Unit
Given $\bm{x}=\{\bm{x}_{i}\}_{i=0}^{T}$, the bidirectional flow-guided propagation unit computes two sequence of hidden states $\{\bm{h}^{\rightarrow}_{i}\}_{i=0}^{T}$ and $\{\bm{h}^{\leftarrow}_{i}\}_{i=0}^{T}$ to capture long-term spatiotemporal dependencies of previous and future frames. It is worth emphasizing that we first fed the features with priors $\bm{p}$ into a ResNet with $N$ residual blocks for feature extraction, which serve as the original feature list $\bm{h}$. We take estimating the forward hidden states $\{\bm{h}^{\rightarrow}_{i}\}_{i=0}^{T}$ as an example. Initially, we estimate the optical flow between the current and previous frames [chan2021basicvsr]:
@@
where $\mathtt{flow}(\cdot)$ denotes a state-of-the-art optical flow estimator [ranjan2017optical]. $\bm{f}_{i\rightarrow{i-1}}$ is then used to align the hidden state $\bm{h}_{i-1}$ backward:
@@
where $\mathtt{calign}(\cdot)$ denotes alignment based on motion compensation (discussed in the next subsection). Subsequently, the aligned previous hidden state $\bm{h}_{i-1\rightarrow i}$ and the current frame $\bm{x}_{i}$ are concatenated along the channel dimension and processed through a ResNet with $N$ residual blocks to compute $\bm{h}^{\rightarrow}_{i}$. The backward hidden state $\bm{h}^{\leftarrow}_{i}$ can be computed by reversing the input sequence and applying the aforementioned formulas. It also needs to be concatenated with the previously computed hidden state and fed into a ResNet with $N$ residual blocks. As prior studies [chan2022basicvsr++] have demonstrated, multiple bidirectional information interaction refines features by propagating intermediate features forward and backward in an alternating manner over time. This allows information from different frames to be revisited for feature refinement. Hence, we employ a two-iteration bidirectional RNN propagation to compute bidirectional hidden states. Specifically, after obtaining forward and backward hidden states via the aforementioned operations, we repeat the process and substitute the original hidden states with the refined ones. To further enhance propagation robustness, we adopt second-order connections, which aggregate information from more spatiotemporal locations, improving the robustness and effectiveness of occluded and fine-detailed regions. More details are provided in the next section. The bidirectional flow-guided propagation unit allows the proposed BasicAVSR to incorporate long-term spatiotemporal context while being flow-aware.